# Tiled Flash Linear Attention: More Efficient Linear RNN and xLSTM Kernels

### v2

@@ Abstract
Linear RNNs with gating recently demonstrated competitive performance compared to Transformers in language modeling. Although their linear compute scaling in sequence length offers theoretical runtime advantages over Transformers, realizing these benefits in practice requires optimized custom kernels, as Transformers rely on the highly efficient Flash Attention kernels (Dao, 2024). Leveraging the chunkwise-parallel formulation of linear RNNs, Flash Linear Attention (FLA) (Yang & Zhang, 2024) shows that linear RNN kernels are faster than Flash Attention, by parallelizing over chunks of the input sequence. However, since the chunk size of FLA is limited, many intermediate states must be materialized in GPU memory. This leads to low arithmetic intensity and causes high memory consumption and IO cost, especially for long-context pre-training. In this work, we present Tiled Flash Linear Attention (TFLA), a novel kernel algorithm for linear RNNs, that enables arbitrary large chunk sizes and high arithmetic intensity by introducing an additional level of sequence parallelization within each chunk. First, we apply TFLA to the xLSTM with matrix memory, the mLSTM (Beck et al., 2024). Second, we propose an mLSTM variant with sigmoid input gate and reduced computation for even faster kernel runtimes at equal language modeling performance. In our speed benchmarks, we show that our new mLSTM kernels based on TFLA outperform highly optimized Flash Attention, Linear Attention and Mamba kernels, setting a new state of the art for efficient long-context sequence modeling primitives.
Our code is available at: https://github.com/NX-AI/mlstm_kernels
Machine Learning, ICML, TiledFlashLinearAttention, TFLA, FLA, FlashLinearAttention, FlashAttention xLSTM, mLSTM, LSTM
@@
1 Introduction
With the trend of training models of ever increasing size with large datasets on thousands of GPUs, it becomes increasingly important to optimize the model architecture as well as its low level implementations for modern hardware. Transformers (Vaswani et al., 2017), which are the core architecture of nowadays state-of-the-art models are highly optimized, but the computational requirements of self-attention scale quadratically with sequence length. This creates significant challenges for both training and inference on long context.
Recently, recurrent alternatives with linear scaling in sequence length (Beck et al., 2024; Sun et al., 2023; Dao & Gu, 2024; Yang et al., 2024b) promise efficiency gains, especially on long sequences and during inference while providing competitive performance. The success of these emerging recurrent architectures is based on two main pillars: (1) a parallel or chunkwise-parallel formulation, which is used in training mode when the full sequence is available beforehand instead of the recurrent formulation and (2) kernel implementations that are close to or exceed training speeds of FlashAttention (Dao, 2024).
Besides the standard recurrent execution, linear RNNs allow for a parallel formulation, which, like Attention, calculates all outputs in parallel. The parallel formulation leverages the insight from linear Attention (Katharopoulos et al., 2020), which showed that kernelized dot-product-based attention can be reinterpreted as a linear RNN with matrix-valued states. Due to the linear nature of the recurrence, the computation can be split into a recurrent part, which computes intermediate RNN states, and a parallel part, which fully utilizes the hardware for computing the outputs in between these intermediate states (Sun et al., 2023; Hua et al., 2022).
Yang et al. (2024b) show that their custom FlashLinearAttention (FLA) kernels based on the chunkwise-parallel formulation of linear RNNs provide faster runtimes than FlashAttention. This is achieved by first splitting the sequence into chunks and materializing the first RNN state of each chunk in GPU memory. Subsequently, in the parallel part they employ one level of sequence parallelism and compute the outputs for each chunk in parallel. For a small chunk size and long sequences, this leads to a large amount of intermediate states to be stored and loaded from GPU memory, which increases memory consumptionand decreases arithmetic intensity. Since modern GPUs see a faster increase in computation throughput than memory bandwidth (Gholami et al., 2024), it is essential to minimize large memory IO and increase arithmetic intensity. A simple approach would be to increase the chunk size. However, the chunk size of FLA is limited by the physical SRAM available on the GPU.
@@ 1 Introduction
In this paper, we implement our TiledFlashLinearAttention algorithm for the xLSTM with matrix memory – the mLSTM (Beck et al., 2024). The mLSTM is a linear RNN that uses exponential gating with scalar gates per head, along with an additional normalizer state for output normalization. This gating mechanism has demonstrated competitive performance compared to Transformers and Mamba on language modeling tasks at moderate scales. However, for comparisons at even larger scales, efficient kernels that leverage the chunkwise-parallel formulation for the mLSTM were still missing. In our speed benchmarks, we show that our new mLSTM kernels based on TFLA outperform highly optimized Attention, Linear Attention and Mamba kernels.
@@
To summarize, in this work, we make the following contributions: (1) We introduce TiledFlashLinearAttention, a new chunkwise-parallel kernel algorithm for Linear RNNs with two levels of sequence parallelism, that enables arbitrary large chunk sizes. (2) We introduce mLSTMsig, a faster mLSTM variant with sigmoid input gate with no performance losses up to 1.4B parameter scales. (3) We improve the training stability and performance of the mLSTM through careful gate initialization guided by our empirical transfer behavior analysis.
@@ 2.1 Recurrent Formulation
${\bm{C}_{t}}$ $=\ {\mathrm{f}_{t}}\ {\bm{C}_{t-1}}\ +\ {\ \mathrm{i}_{t}}\ {\bm{k}_{t}\ \bm{v}_{t}^{\top}}$ (2)
${\bm{n}_{t}\ \!}$ $=\ {\mathrm{f}_{t}}\ {\bm{n}_{t-1}\ \!}\ +\ {\ \mathrm{i}_{t}}\ {\bm{k}_{t}}$ (3)
@@
$\bm{h}_{t}$ $=\ {\mathbf{o}_{t}}\ \odot\ \mathrm{NORM}(\,\widetilde{\bm{h}}_{t}\,)$ (5)
The gate activations are computed as:
${\mathrm{f}_{t}}$ $={\exp\left(\log\sigma(\tilde{\mathrm{f}}_{t})+m_{t-1}-m_{t}\right)\,}$ (6)
${\ \mathrm{i}_{t}}$ $={\exp(\tilde{\mathrm{i}}_{t}-m_{t})}$ (7)
${\mathbf{o}_{t}}$ $=\ \sigma\left(\tilde{\mathbf{o}}_{t}\right)$ (8)
The query, key, and value vectors $\bm{q}_{t},\bm{k}_{t}\in\mathbb{R}^{d_{qk}},\,\bm{v}_{t}\in\mathbb{R}^{d_{hv}}$ are computed as $\{\bm{q}_{t},\bm{k}_{t},\bm{v}_{t}\}=\ \bm{W}_{\{q,k,v\}}\ \bm{x}_{t}\ +\ \bm{b}_{\{q,k,v\}}$. The scalar input and forget gates $\mathrm{i}_{t},\mathrm{f}_{t}\in\mathbb{R}$ are computed from the pre-activations $\{\tilde{\mathrm{i}}_{t},\tilde{\mathrm{f}}_{t}\}=\ \bm{w}^{\top}_{\{\mathrm{i},\mathrm{f}\}}\ \bm{x}_{t}\ +\ b_{\{\mathrm{i},\mathrm{f}\}}$ and the vector output gate $\mathbf{o}_{t}\in\mathbb{R}^{d_{hv}}$ is computed from the pre-activation $\tilde{\mathbf{o}}_{t}\ =\ \bm{W}_{\mathbf{o}}\ \bm{x}_{t}\ +\ \bm{b}_{\mathbf{o}}$ with the sigmoid function $\sigma$. The norm layer $\mathrm{NORM}$ in (5) can be either RMS norm (Zhang & Sennrich, 2019) or LayerNorm (Ba et al., 2016). Typically, multiple of these cells operate simultaneously as parallel heads, similar to Transformers (Vaswani et al., 2017).
@@ 2.2 Chunkwise-Parallel Formulation
Given the logarithmic forget gates $\bar{\bm{\mathrm{f}}}^{(k)}=\log\sigma(\tilde{\bm{\mathrm{f}}}^{(k)})\in\mathbb{R}^{L}$ and input gates $\bar{\bm{\mathrm{i}}}^{(k)}=\log\exp(\tilde{\bm{\mathrm{i}}}^{(k)})\in\mathbb{R}^{L}$, we can compute the logarithmic chunkwise gates as [^1] See Appendix B.2 for more details on the chunkwise gates.
$\mathrm{g}_{k}$ $=\mathrm{sum}\left(\bar{\bm{\mathrm{f}}}^{(k)}\right)\quad\in\mathbb{R},$ (9)
$\bm{\mathrm{b}}_{k}$ $=\mathrm{cumsum}\left(\bar{\bm{\mathrm{f}}}^{(k)}\right)\;\in\mathbb{R}^{L},$ (10)
$\bm{\mathrm{a}}_{k}$ $=\mathrm{rev\_cumsum}\left(\bar{\bm{\mathrm{f}}}^{(k)}\right)+\bar{\bm{\mathrm{i}}}^{(k)}\;\in\mathbb{R}^{L}.$ (11)
@@
In Figure 2, we show the summed forget gates $\mathrm{g}_{k}$ contain the forget gate contribution of all forget gates within a chunk. The cumulative forget gate vectors $\bm{\mathrm{b}}_{k}$ contain the forget gate contributions from the beginning of the chunk up to the current time step within the current chunk. The cumulative input gate vectors $\bm{\mathrm{a}}_{k}$ contain the input gates for every timestep as well as the forget gate contributions from the current time step to the end of the chunk.
@@
where $\bar{\mathrm{g}}_{k}$ and $\overline{\bm{\mathrm{a}}}_{k}$ are the stabilized chunkwise gates.11footnotemark: 1 This recurrent part resembles the fully recurrent formulation in Section 2.1, but instead of computing the intermediate states for every timestep $t$, we compute them directly for every $L$ time steps without materializing the states in between.
@@
$\widetilde{\bm{\mathrm{D}}}^{(k)}$ =$=\begin{cases}-\infty&\text{for}\;i<j\\ \bm{\mathrm{b}}_{k}-\bm{\mathrm{b}}_{k}^{\top}+{\overline{\bm{\mathrm{i}}}^{(k)}}^{\top}&\text{for}\;i\geqslant j\end{cases}$ (14)
$\bm{\mathrm{S}}^{(k)}$ $=\frac{1}{\sqrt{d_{qk}}}{\bm{Q}^{(k)}}{\bm{K}^{(k)}}^{\top}$ (15)
$\overline{\bm{\mathrm{S}}}^{(k)}$ $=\bm{\mathrm{S}}^{(k)}\odot\bm{\mathrm{D}}^{(k)},$ (16)
@@
$\bm{\mathrm{H}}_{\text{inter}}^{(k)}$ $=\left(\overline{\bm{\mathrm{b}}}_{k}\odot\frac{\bm{Q}^{(k)}}{\sqrt{d_{qk}}}\right)\bm{C}_{k-1}=\overline{\bm{Q}}^{(k)}\bm{C}_{k-1}$ (17)
@@
where $\bm{\mathrm{h}}^{(k)}_{\text{denom}}\in\mathbb{R}^{L}$ is a normalization factor.
Appendix B.2 and B.3 provide a detailed description of the chunkwise-parallel forward and backward pass. Appendix F provides the FLOP and memory operation counts for all formulations.
@@ 3 Tiled Flash Linear Attention
FlashLinearAttention (Yang et al., 2024b) introduces a fast kernel algorithm for the chunkwise formulation for Linear Attention (cf. Section 2.2 without gates) and shows that their implementation is faster than optimized FlashAttention (Dao, 2024). This speedup is achieved by single level sequence parallelism, where the states $\bm{C}_{k}$ are first materialized in GPU memory and then the outputs $\bm{\mathrm{H}}^{(k)}$ are computed in parallel. However, since in FlashLinearAttention the chunk size parameter determines the tile sizes in SRAM, the maximum chunk size is limited (typically $L=64$) by the physical SRAM size of the GPU. Therefore, we have to materialize many states in HBM, where the number of states is $N_{c}=\lceil T/L\rceil$. This leads to low arithmetic intensity and high GPU memory consumption, which poses challenges especially for long-context pre-training. To address this issue, we introduce TiledFlashLinearAttention (TFLA) in Section 3.2, which enables arbitrary large chunk sizes.
We begin with a review on the fundamentals of writing efficient kernels in Section 3.1. Since we perform our experiments on NVIDIA GPUs, our review is targeted towards NVIDIA’s terminology, though the principles also apply to other hardware. For a more extensive overview we refer to (Spector et al., 2024). Finally, in Appendix G, we analyze the theoretical runtime of our TFLA kernels.
3.1 GPU Fundamentals
3.2 More Efficient Kernels via Two Level Sequence Parallelism
TiledFlashLinearAttention (TFLA) enables fast kernels and a trade-off between memory consumption and computational efficiency by introducing two levels of sequence parallelism (see Figure 6). The first level is the parallelization over the chunks of the sequence, which requires to compute and materialize intermediate states $\bm{C}_{k}$ in GPU HBM. For this we use a recurrent kernel similar to previous work (Yang et al., 2024b). The second level is the parallelization within each chunk, which is achieved by tiling the intra chunk attention matrix along the chunk dimension. This second level of parallelism enables large chunk sizes and hence reduces the memory consumption for the intermediate states as we have to store and load $N_{c}=\lceil T/L\rceil$ intermediate states in HBM on each kernel call, where $T$ is the sequence length and $L$ is the chunk size. In addition to the two levels of sequence parallelism and the naive parallelization over the batch and head dimensions, TFLA also parallelizes over the embedding dimension. This enables arbitrary large head dimensions and results in a massive parallelization over five dimensions, which is crucial for achieving high performance on modern GPUs.
@@ 3.2 More Efficient Kernels via Two Level Sequence Parallelism
We review the matrix multiplication operations of the intra-chunk parallel part of the mLSTM in order to show how we efficiently parallelize these operations. For simplicity we omit the the gate computations and normalization, as these do not influence the work partitioning. We also omit the leading batch, head and chunk dimension, over which we can parallelize naively as they do not interact with the matrix multiplication (see Table 1).
Figure 3: TFLA Forward Pass Tiling. We loop over $B_{Lkv}$ and $B_{dqk}$ (indicated by arrows) and parallelize over $B_{Lhq}$ and $B_{dhv}$ (indicated by dashed lines) blocks. $\bigoplus$ denotes block-wise accumulation.
In simplified form, the intra-chunk parallel forward pass of the mLSTM (and other linear RNNs) for a chunk $k$ can be written as three matrix multiplications, which we fuse into a single kernel:
$\underset{(L_{hq}\times d_{hv})}{\bm{\mathrm{H}}^{(k)}}$ $=\underbrace{\left(\underset{(L_{hq}\times d_{qk})}{\bm{Q}^{(k)}}\ \underset{(d_{qk}\times L_{kv})}{{\bm{K}^{(k)}}^{\top}}\right)\ \underset{(L_{kv}\times d_{hv})}{\bm{V}^{(k)}}}_{\bm{\mathrm{H}}^{(k)}_{\text{intra}}}$ (20)
$+\underbrace{\underset{(L_{hq}\times d_{qk})}{\bm{Q}^{(k)}}\ \underset{(d_{qk}\times d_{hv})}{\bm{C}_{k-1}}}_{\bm{\mathrm{H}}^{(k)}_{\text{inter}}}$
@@
Kernel $L_{hq}$ $L_{kv}$ $d_{qk}$ $d_{hv}$ Thread Block Grid
$\bm{\mathrm{H}}^{(k)}$ P L L P $\left(\frac{d_{hv}}{B_{dhv}},\frac{L_{hq}}{B_{Lhq}},\dots\right)$
$\delta\bm{Q}^{(k)}$ P L P L $\left(\frac{d_{qk}}{B_{dqk}},\frac{L_{hq}}{B_{Lhq}},\dots\right)$
$\delta\bm{K}^{(k)}$ L P P L $\left(\frac{d_{qk}}{B_{dqk}},\frac{L_{kv}}{B_{Lkv}},\dots\right)$
$\delta\bm{V}^{(k)}$ L P L P $\left(\frac{d_{hv}}{B_{dhv}},\frac{L_{kv}}{B_{Lkv}},\dots\right)$

### v3

@@ Abstract
Linear RNNs with gating recently demonstrated competitive performance compared to Transformers in language modeling. Although their linear compute scaling in sequence length offers theoretical runtime advantages over Transformers, realizing these benefits in practice requires optimized custom kernels, as Transformers rely on the highly efficient Flash Attention kernels (Dao, 2024). Leveraging the chunkwise-parallel formulation of linear RNNs, Flash Linear Attention (FLA) (Yang & Zhang, 2024) shows that linear RNN kernels are faster than Flash Attention, by parallelizing over chunks of the input sequence. However, since the chunk size of FLA is limited, many intermediate states must be materialized in GPU memory. This leads to low arithmetic intensity and causes high memory consumption and IO cost, especially for long-context pre-training. In this work, we present Tiled Flash Linear Attention (TFLA), a novel kernel algorithm for linear RNNs, that enables arbitrary large chunk sizes and high arithmetic intensity by introducing an additional level of sequence parallelization within each chunk. First, we apply TFLA to the xLSTM with matrix memory, the mLSTM (Beck et al., 2024). Second, we propose an mLSTM variant with sigmoid input gate and reduced computation for even faster kernel runtimes at equal language modeling performance. In our speed benchmarks, we show that our new mLSTM kernels based on TFLA outperform highly optimized Flash Attention, Linear Attention and Mamba kernels, setting a new state of the art for efficient long-context sequence modeling primitives. Our code is available at: https://github.com/NX-AI/mlstm_kernels
1 Introduction
With the trend of training models of ever increasing size with large datasets on thousands of GPUs, it becomes increasingly important to optimize the model architecture as well as its low-level implementations for modern hardware. Transformers (Vaswani et al., 2017), which are the core architecture of nowadays state-of-the-art models are highly optimized, but the computational requirements of self-attention scale quadratically with sequence length. This creates significant challenges for both training and inference on long context.
Recently, recurrent alternatives with linear scaling in sequence length (Beck et al., 2024; Sun et al., 2023; Dao & Gu, 2024; Yang et al., 2024b) promise efficiency gains, especially on long sequences and during inference while providing competitive performance. The success of these emerging recurrent architectures is based on two main pillars: (1) A parallel or chunkwise-parallel formulation (Sun et al., 2023; Hua et al., 2022), which, like Attention, calculates all outputs in parallel during training, and (2) kernel implementations that are close to or exceed training speeds of FlashAttention (Dao, 2024).
Yang et al. (2024b) show that their custom Flash Linear Attention (FLA) kernels, based on the chunkwise-parallel formulation of linear RNNs, achieve faster runtimes than FlashAttention.
@@ 1 Introduction
They accomplish this by dividing the sequence into chunks and recurrently materializing only the initial RNN state of each chunk in GPU memory. Subsequently, in the parallel part they employ one level of sequence parallelism and compute the outputs for each chunk in parallel. For a small chunk size and long sequences, this leads to a large amount of intermediate states to be stored and loaded from GPU memory, which increases memory consumptionand decreases arithmetic intensity. Since modern GPUs see a faster increase in computation throughput than memory bandwidth (Gholami et al., 2024), it is essential to minimize large memory IO and increase arithmetic intensity. A simple approach would be to increase the chunk size. However, the chunk size of FLA is limited by the physical SRAM available on the GPU.
@@
In this paper, we implement our Tiled Flash Linear Attention algorithm for the xLSTM with matrix memory – the mLSTM Beck et al. (2024). The mLSTM is a linear RNN that uses exponential gating with scalar gates per head, along with an additional normalizer state for output normalization. This gating mechanism has demonstrated competitive performance compared to Transformers and Mamba on language modeling tasks at moderate scales. However, for comparisons at even larger scales, efficient kernels that leverage the chunkwise-parallel formulation for the mLSTM were still missing. In our speed benchmarks, we show that our new mLSTM kernels based on TFLA outperform highly optimized Attention, Linear Attention and Mamba kernels.
@@
To summarize, in this work, we make the following contributions: (1) We introduce Tiled Flash Linear Attention, a new chunkwise-parallel kernel algorithm for Linear RNNs with two levels of sequence parallelism, that enables arbitrary large chunk sizes and apply it to the mLSTM (Beck et al., 2024). (2) We introduce mLSTMsig, a faster mLSTM variant with sigmoid input gate with no performance losses up to 1.4B parameter scales. (3) We improve the training stability and performance of the mLSTM through careful gate initialization guided by our empirical transfer behavior analysis.
@@ 2.1 Recurrent Formulation
${\bm{C}_{t}}\$$=\ {\mathrm{f}_{t}}\ {\bm{C}_{t-1}}\ +\ {\ \mathrm{i}_{t}}\ {\bm{k}_{t}\ \bm{v}_{t}^{\top}}$ (2)
${\bm{n}_{t}\ \!}\$$=\ {\mathrm{f}_{t}}\ {\bm{n}_{t-1}\ \!}\ +\ {\ \mathrm{i}_{t}}\ {\bm{k}_{t}}$ (3)
@@
$\bm{h}_{t}\$$=\ {\mathbf{o}_{t}}\ \odot\ \mathrm{NORM}(\,\widetilde{\bm{h}}_{t}\,)$ (5)
[…]
@@ 2.2 Chunkwise-Parallel Formulation
Given the logarithmic forget gates $\bar{\bm{\mathrm{f}}}^{(k)}=\log\sigma(\tilde{\bm{\mathrm{f}}}^{(k)})\in\mathbb{R}^{L}$ and input gates $\bar{\bm{\mathrm{i}}}^{(k)}=\log\exp(\tilde{\bm{\mathrm{i}}}^{(k)})\in\mathbb{R}^{L}$, we can compute the logarithmic chunkwise gates $\mathrm{g}_{k}\in\mathbb{R}$, and $\bm{\mathrm{b}}_{k},\bm{\mathrm{a}}_{k}\in\mathbb{R}^{L}$ as $\mathrm{g}_{k}=\mathrm{sum}\left(\bar{\bm{\mathrm{f}}}^{(k)}\right)$, $\bm{\mathrm{b}}_{k}=\mathrm{cumsum}\left(\bar{\bm{\mathrm{f}}}^{(k)}\right)$, and $\bm{\mathrm{a}}_{k}=\mathrm{rev\_cumsum}\left(\bar{\bm{\mathrm{f}}}^{(k)}\right)+\bar{\bm{\mathrm{i}}}^{(k)}$. We refer to Appendix B.2 for more details on the chunkwise gates. In Figure 2, we show the summed forget gates $\mathrm{g}_{k}$ contain the forget gate contribution of all forget gates within a chunk. The cumulative forget gate vectors $\bm{\mathrm{b}}_{k}$ contain the forget gate contributions from the beginning of the chunk up to the current time step within the current chunk. The cumulative input gate vectors $\bm{\mathrm{a}}_{k}$ contain the input gates for every timestep as well as the forget gate contributions from the current time step to the end of the chunk.
@@
where $\bar{\mathrm{g}}_{k}$ and $\overline{\bm{\mathrm{a}}}_{k}$ are the stabilized chunkwise gates.This recurrent part resembles the fully recurrent formulation in Section 2.1, but instead of computing the intermediate states for every timestep $t$, we compute them directly for every $L$ time steps without materializing the states in between.
@@
$\widetilde{\bm{\mathrm{D}}}^{(k)}$ $=\begin{cases}-\infty&\text{for}\;i<j\\ \bm{\mathrm{b}}_{k}-\bm{\mathrm{b}}_{k}^{\top}+{\overline{\bm{\mathrm{i}}}^{(k)}}^{\top}&\text{for}\;i\geqslant j\end{cases}$ (8)
$\overline{\bm{\mathrm{S}}}^{(k)}$ $=\left(\frac{1}{\sqrt{d_{qk}}}{\bm{Q}^{(k)}}{\bm{K}^{(k)}}^{\top}\right)\odot\bm{\mathrm{D}}^{(k)}=\bm{\mathrm{S}}^{(k)}\odot\bm{\mathrm{D}}^{(k)},$ (9)
@@
$\bm{\mathrm{H}}_{\text{inter}}^{(k)}=\left(\overline{\bm{\mathrm{b}}}_{k}\odot\frac{\bm{Q}^{(k)}}{\sqrt{d_{qk}}}\right)\bm{C}_{k-1}=\overline{\bm{Q}}^{(k)}\bm{C}_{k-1},\quad\quad\bm{\mathrm{H}}_{\text{intra}}^{(k)}=\overline{\bm{\mathrm{S}}}^{(k)}\bm{V}^{(k)},$ (10)
@@
where $\bm{\mathrm{h}}^{(k)}_{\text{denom}}\in\mathbb{R}^{L}$ is a normalization factor. Appendix B.2 and B.3 provide a detailed description of the chunkwise-parallel forward and backward pass. Appendix F provides the FLOP and memory operation counts for all formulations.
@@ 3 Tiled Flash Linear Attention
Flash Linear Attention (Yang et al., 2024b) introduces a fast kernel algorithm for the chunkwise formulation for Linear Attention (cf. Section 2.2 without gates) and shows that their implementation is faster than optimized FlashAttention (Dao, 2024). This speedup is achieved by single level sequence parallelism, where the states $\bm{C}_{k}$ are first materialized in GPU memory and then the outputs $\bm{\mathrm{H}}^{(k)}$ are computed in parallel. However, since in Flash Linear Attention the chunk size parameter determines the tile sizes in SRAM, the maximum chunk size is limited (typically $L=64$) by the physical SRAM size of the GPU. Therefore, we have to materialize many states in HBM, where the number of states is $N_{c}=\lceil T/L\rceil$. This leads to low arithmetic intensity and high GPU memory consumption, which poses challenges especially for long-context pre-training.
We begin with a brief review of fundamentals on GPUs for writing efficient kernels in Appendix C.1.
More Efficient Kernels via Two Level Sequence Parallelism.
To address the issue of limited chunk sizes, Tiled Flash Linear Attention (TFLA) introduces two levels of sequence parallelism, which enables fast kernels and a trade-off between memory consumption and computational efficiency (see Figure 6). The first level is the parallelization over the chunks of the sequence, which requires to compute and materialize intermediate states $\bm{C}_{k}$ in GPU High Bandwidth memory (HBM). For this we use a recurrent kernel similar to previous work (Yang et al., 2024b). The second level is the parallelization within each chunk, which is achieved by tiling the intra chunk attention matrix along the chunk dimension. This second level of parallelism enables large chunk sizes and hence reduces the memory consumption for the intermediate states as we have to store and load $N_{c}=\lceil T/L\rceil$ intermediate states in HBM on each kernel call, where $T$ is the sequence length and $L$ is the chunk size. In addition to the two levels of sequence parallelism and the naive parallelization over the batch and head dimensions, TFLA also parallelizes over the embedding dimension. This enables arbitrary large head dimensions and results in a massive parallelization over five dimensions, which is crucial for achieving high performance on modern GPUs. We analyze the theoretical runtime of our TFLA kernels in Appendix G.
@@
We review the matrix multiplication operations of the intra-chunk parallel part of the mLSTM in order to show how we efficiently parallelize these operations. For simplicity we omit the the gate computations and normalization, as these do not influence the work partitioning. We also omit the leading batch, head and chunk dimension, over which we can parallelize naively as they do not interact with the matrix multiplication (see Table 1). In simplified form, the intra-chunk parallel forward pass of the mLSTM (and other linear RNNs) for a chunk $k$ can be written as three matrix multiplications, which we fuse into a single kernel:
$\underset{(L_{hq}\times d_{hv})}{\bm{\mathrm{H}}^{(k)}}$ $=\underbrace{\left(\underset{(L_{hq}\times d_{qk})}{\bm{Q}^{(k)}}\ \underset{(d_{qk}\times L_{kv})}{{\bm{K}^{(k)}}^{\top}}\right)\ \underset{(L_{kv}\times d_{hv})}{\bm{V}^{(k)}}}_{\bm{\mathrm{H}}^{(k)}_{\text{intra}}}+\underbrace{\underset{(L_{hq}\times d_{qk})}{\bm{Q}^{(k)}}\ \underset{(d_{qk}\times d_{hv})}{\bm{C}_{k-1}}}_{\bm{\mathrm{H}}^{(k)}_{\text{inter}}}$ (12)
Figure 3: TFLA Intra-Chunk Tiling. We loop over $B_{Lkv}$ and $B_{dqk}$ (indicated by arrows) and parallelize over $B_{Lhq}$ and $B_{dhv}$ (indicated by dashed lines) blocks. $\bigoplus$ denotes block-wise accumulation.
@@
Kernel $L_{hq}$ $L_{kv}$ $d_{qk}$ $d_{hv}$ Thread Block Grid $\bm{\mathrm{H}}^{(k)}$ P L L P $\left(\frac{d_{hv}}{B_{dhv}},\frac{L_{hq}}{B_{Lhq}},\dots\right)$ $\delta\bm{Q}^{(k)}$ P L P L $\left(\frac{d_{qk}}{B_{dqk}},\frac{L_{hq}}{B_{Lhq}},\dots\right)$ $\delta\bm{K}^{(k)}$ L P P L $\left(\frac{d_{qk}}{B_{dqk}},\frac{L_{kv}}{B_{Lkv}},\dots\right)$ $\delta\bm{V}^{(k)}$ L P L P $\left(\frac{d_{hv}}{B_{dhv}},\frac{L_{kv}}{B_{Lkv}},\dots\right)$